OpenStudy (anonymous):

In 3.45 mg of boron there are how many of the following? grams of boron? moles of boron? boron atoms?

OpenStudy (anonymous):

\[3.45mg \times \frac{1g}{1000mg}=0.00345g\]This is basic dimensional anaysis. You will have better luck on the chemistry site for the other two questions, I think.

OpenStudy (anonymous):

moles of boron = .00345/relative atomic mass of boron in grams for the third part you need to use the number of atoms in 1 mole of born which is Avogadros Number whose value i cant recall

OpenStudy (chaise):

Moles = Mass(g)/MW Moles = 0.003.45/10.811 Moles = 0.32 There are 0.00032 moles of Boron in 3.45mg or 0.00345g.

OpenStudy (anonymous):

6.02x10 to 23rd power

OpenStudy (anonymous):

thanks all of you so much!!!

OpenStudy (anonymous):

yw